# Construction of Exponential Families from Statistical Manifolds

### v1

@@ Abstract.
This paper addresses the fundamental question of constructing exponential families from statistical manifolds. We show that every compact statistical manifold admits a singular foliation whose leaves are Hessian manifolds. Moreover, any non-flat, compact, orientable 3-dimensional leaf appears as a quotient of an exponential family and has only odd Betti.
@@ 1. Introduction
Information geometry is a rapidly developing discipline that establishes a profound connection between differential geometry and statistical theory through the Fisher information metric (see [amari2000methods, amari2016information, lauritzen1987statistical, ay2017information]). Numerous researchers have explored the construction of meaningful distances between probability distributions, acknowledging its critical role in diverse areas including theoretical statistics, information theory, signal processing, quantitative finance, and machine learning. To solve this problem, Rao [rao1945information] introduced the geodesic distance derived from the Fisher information metric, now universally recognized as the Fisher-Rao distance (or Rao’s distance). The Fisher metric, also called the information metric, is generally positive semi-definite but can be degenerate. Cencov ([cencov2000statistical], p. 156) proved the uniqueness result for Fisher’s information metric on finite sample spaces. Subsequently, in [ay2015information], the authors extended Cencov’s result to infinite sample spaces. The Fisher information metric is invariant under parametrization on a sample space $\Xi$ and covariant under parametrization on a parametric space $\Theta$, see [calin2014geometric]. The Fisher-Rao distance is inherently designed to be invariant under diffeomorphisms of both the sample space $\Xi$ and the parameter space $\Theta$. The Fisher metric has been widely used in many other areas of geometry (see [hitchin2006geometry, fujiwara2020dually, groisser1997instantons, ciaglia2021focus, murray1993information, matsueda2017banados, naudts2009q, petz2007quantum, lesniewski1999monotone, facchi2010classical, barbaresco2014koszul]).
Information geometry exhibits deep connections with several mathematical disciplines, including Hessian geometry, Kähler geometry, complex geometry, special Kähler geometry, and complex analysis. One of the most fundamental problems in information geometry is the realization of a statistical manifold as a statistical model, that is, a manifold equipped with dual connections with respect to a Fisher metric. This question was resolved by Hông Vân Lê [le2006statistical]. Another, much more challenging problem is to determine which classes of statistical manifolds can be realized as exponential families; this problem remains open. Despite its rapid development, information geometry still contains many significant open problems [sei2025open].
This paper is organized as follows. Section 2 reviews the relevant background notions in information geometry. Section 3 is devoted to the proof of the main result of the paper.
2. Basics Notions
@@ 2.1. Statistical models
[amari1987differential, amari2000methods, amari2012differential, amari2014curvature, amari2016information, ay2017information, calin2014geometric] Let $\mathcal{S}$ be a family of probability distributions. Assume that $\mathcal{S}$ is represented as
@@
where $\Theta$ is a domain in $\mathbb{R}^{n}$ with Cartesian coordinates $(\theta^{1},\theta^{2},\dots,\theta^{n})$ (called the parameter space). In addition, we assume that the mapping $\theta\mapsto p(\cdot,\theta)$ is bijective. Then $\mathcal{S}$ is said to be a statistical model of probability distributions determined by $p$.
@@
$\Gamma^{(\alpha)}_{ijk}(\theta)=\mathbb{E}_{p(\cdot,\theta)}\left[\left(\frac{\partial^{2}\log p(x,\theta)}{\partial\theta^{i}\partial\theta^{j}}+\frac{1-\alpha}{2}\frac{\partial\log p(x,\theta)}{\partial\theta^{i}}\frac{\partial\log p(x,\theta)}{\partial\theta^{j}}\right)\log p(x,\theta)\cdot\frac{\partial\log p(x,\theta)}{\partial\theta^{k}}\right].$
In the literature, the family of $\alpha$-connections $\nabla^{(\alpha)}$ is commonly referred to as the Amari-Chentsov $\alpha$-connections or Amari–Čencov $\alpha$-connections [amari2000methods] . This terminology honors the foundational contributions of Shun-ichi Amari [amari1987differential, amari2012differential] and Nikolai Chentsov[chentsov1982statistical] to information geometry. The Amari-Chentsov $\alpha$-connections provide a unified framework for studying the differential geometric structure of statistical models, with the Fisher metric as the central Riemannian structure and the $\alpha$-connections forming a dualistic structure on the statistical manifold.
The $\alpha$-connection $\nabla^{(\alpha)}$ satisfies the following fundamental properties:
@@
This corresponds exactly to the Gaussian distribution $\mathcal{N}(\mu,\sigma)$.
For an exponential family with natural parameters $\theta=(\theta^{1},\dots,\theta^{n})$, the Fisher metric and the cubic form are given in terms of the cumulant generating function $\psi(\theta)$ by:
@@ 2.2. Statistical manifold
Building upon the geometric insights of Amari and Chentsov, Steffen L. Lauritzen provided a formalize the notion of a statistical manifold.
@@
The following construction provides a fundamental example of how statistical structures naturally arise on Riemannian manifolds equipped with a distinguished vector field.
Let $(M,g)$ be a Riemannian manifold with Levi–Civita connection $\nabla^{\mathrm{LC}}$, and let $\xi\in TM\setminus\{0\}$ be a non-vanishing vector field. For all $X,Y\in\mathfrak{X}(M)$, define
@@
Then $(M,g,\nabla,\nabla^{*})$ is a statistical manifold.
Starting from our previous construction, on an almost contact metric manifold $(M^{2d+1},\phi,g,\alpha,\xi)$ with $\nabla^{LC}$ the Levi–Civita connection of the metric $g$ (see [blair2010riemannian] for exposition on almost contact metric geometry), we can construct the following one-parameter families of statistical connections:
@@
For a statistical manifold $(M,g,\nabla,\nabla^{*})$. The connection $\nabla$ is flat if and only if $\nabla^{*}$ is flat. In this case, we say that $(M,g,\nabla,\nabla^{*})$ is a dually flat manifold ([amari2016information]).
It was shown in [ay2002dually] that any compact dually flat manifold necessarily has an infinite fundamental group. Consequently, the sphere $\mathbb{S}^{q}$, for $q>1$, cannot be a dually flat manifold.
@@ 2.3. Hessian manifold
Hessian manifolds exhibit particularly rich topological and geometric structures and share a deep relationship with Kähler manifolds. This relationship is especially manifested in the fact that the tangent bundle of any Hessian manifold is naturally endowed with a Kähler metric induced by the Hessian structure itself [shima2007geometry, dombrowski1962geometry, satoh2019almost]. Such manifolds occupy a central place in information geometry, Kähler geometry, special Kähler geometry, and complex analysis. They also play an important role in various applied fields, including convex optimization, statistics, combinatorics, Souriau’s thermodynamics on Lie groups, and quantitative finance.
@@
where $(x_{1},\dots,x_{n})$ is a system of affine coordinates with respect to $\nabla$.
A locally flat manifold $(M,\nabla)$ equipped with a Hessian metric $g$ is called a Hessian manifold and is denoted by $(M,g,\nabla)$.
@@
On a orientable Hessian manifold $(M,g,\nabla)$, using the connection $\nabla$ and the volume element induced by $g$, Koszul introduced a closed 1-form, which Shima [shima2007geometry] refers to as the first Koszul form, associated with the Hessian structure $(\nabla,g)$.
@@ 3. Main Results
Theorem 3.1 ([le2006statistical]).
Every smooth compact statistical manifold $(M^{n},g,T)$ of finite dimension admits, for a sufficiently large integer $N$, an isostatistical embedding (i.e., an embedding that preserves the tensor $T$) into the space $\mathrm{Cap}_{N}$ of probability densities on $\Omega_{N}$, where $\Omega$ is a differentiable manifold endowed with a Borel measure.
@@
On any compact statistical manifold, there exists a singular foliation whose leaves are Hessian manifolds. Every compact, orientable leaf is either a finite quotient of a flat torus (flat) or a mapping torus with periodic glueing map. If such a leaf is three-dimensional and non-flat, then all of its Betti numbers are odd, and it arises as a quotient of an exponential family.
@@
The proof proceeds in four steps.
@@ 3. Step 3:
    We thus deduce that $\widetilde{Z}$ is the Reeb vector field associated with the cosymplectic structure $(\widetilde{\beta},\Omega)$ (see [li2008topology] ). Since $\widetilde{Z}$ is a Killing vector field on $L_{\nabla}$, it follows from [bazzoni2015k] that $(L_{\nabla},\widetilde{\beta},\Omega)$ is a K-cosymplectic manifold. By [bazzoni2015k, Proposition 2.8], $L_{\nabla}$ admits an almost contact metric structure $(\widetilde{\beta},\widetilde{Z},\phi,h)$ with $\mathcal{L}_{\widetilde{Z}}h=0$.
@@
    $L_{\nabla}\cong\Gamma\backslash(\mathbb{H}^{2}\times\mathbb{R})\quad\text{with}\quad\Gamma\subset\mathrm{Isom}_{0}(\mathbb{H}^{2}\times\mathbb{R})\cong\mathrm{Isom}_{0}(\mathbb{H}^{2})\times\mathrm{Isom}_{0}(\mathbb{R}).$
@@ 4. Step 4:
    $P((x,y,z),(\xi,\eta,\zeta))=\exp\left(-(x\xi+y\eta+z\zeta)+\frac{3}{2}\log q-\log\chi(0,0,1)\right).$
@@
    Introducing cylindrical coordinates $(t,r,\alpha)$ by
@@
    Here $(r,\alpha)$ are polar coordinates on $\mathbb{H}^{2}$ and $t$ is the fiber coordinate. Thus, $(\mathbb{H}^{2}\times\mathbb{R},g)$ is isometric to $(\Theta,g^{F})$.
    We deduce that, the leaf $L_{\nabla}$ of the associated foliation $\mathcal{F}_{\nabla}$ satisfies
    $L_{\nabla}\cong\Lambda\backslash\Theta,\qquad\Lambda\subset\mathrm{Isom}_{0}(\Theta).$
@@
[…]
    where $\{P_{\theta}:\theta\in\Theta\}$ is an exponential family. This completes the proof.
@@
$L_{\nabla}\cong\Gamma\backslash(\mathbb{H}^{2}\times\mathbb{R}),\qquad\Gamma\subset\mathrm{Isom}_{0}(\mathbb{H}^{2}\times\mathbb{R}),$
is a Seifert manifold with Euler number 0 and hyperbolic orbifold base $B$ (since $\chi_{\mathrm{orb}}(B)<0$). In other words, $L_{\nabla}$ admits a $\mathbb{H}^{2}\times\mathbb{R}$-geometry.
From [martelli2016introduction, Prop. 10.3.26], $L_{\nabla}$ is covered by $\Sigma_{g}\times\mathbb{S}^{1}$ for some $g>1$. Hence
$L_{\nabla}\cong\Pi\backslash(\Sigma_{g}\times\mathbb{S}^{1}),$
where $\Pi$ is a finite group acting freely and properly on $\Sigma_{g}\times\mathbb{S}^{1}$.
Since $\Sigma_{g}$ admits an $\mathbb{RP}^{2}$-structure (see [benzecri1960varietes, p. 243] and [goldman2022geometric, Cor. 6.5.2]), the universal cover of $\Sigma_{g}\times\mathbb{S}^{1}$, namely $\mathbb{H}^{2}\times\mathbb{R}$, develops into the Lorentz cone $\Theta$. Consequently,
$\Sigma_{g}\times\mathbb{S}^{1}\cong(\Gamma\times\langle\lambda\rangle)\backslash\Theta,\qquad\Gamma\subset\mathrm{SO}(2,1),\quad\lambda>1.$
@@
Finally, by using the Koszul–Vinberg characteristic function $\chi$ on the sharp convex cone $\Theta$, and arguing as above, the result follows.

### v2

@@ Abstract.
We investigate the construction of exponential families from statistical manifolds, a central problem in information geometry. We prove that every compact statistical manifold admits a singular foliation whose leaves are Hessian manifolds. In particular, any non-flat, compact, orientable $3$-dimensional leaf arises as a quotient of an exponential family and has only odd Betti numbers. Our approach is constructive: we explicitly describe the foliation and analyze the geometric and topological properties of its leaves. We show that compact orientable leaves are either finite quotients of flat torus or mapping torus with periodic monodromy. In three dimensions, non-flat leaves admit a co-Kähler structure, which allows us to realize them as explicit exponential families parametrized by a Lorentz cone. These results establish a concrete bridge between abstract statistical manifolds and exponential families, highlighting deep connections between information geometry, differential geometry, and the topology of $3$-manifolds.
@@ 1. Introduction
Information geometry is a rapidly expanding field that reveals a deep interplay between differential geometry and statistical theory through the Fisher information metric (see [amari2000methods, molitor2013exponential, amari2016information, lauritzen1987statistical, ay2017information]). Considerable attention has been devoted to the problem of defining meaningful distances between probability distributions, given their fundamental importance in a wide range of applications such as theoretical statistics, information theory, signal processing, quantitative finance, and machine learning. In this context, Rao [rao1945information] introduced the geodesic distance induced by the Fisher information metric, which is now commonly referred to as the Fisher–Rao distance (or Rao’s distance). The Fisher metric, also known as the information metric, is in general positive semi-definite, although it may be degenerate. Čencov proved a fundamental uniqueness theorem for the Fisher information metric on finite sample spaces ([cencov2000statistical], p. 156). This result was later extended to infinite sample spaces in [ay2015information].
The Fisher information metric is invariant under reparametrizations of the sample space $\Xi$ and covariant with respect to reparametrizations of the parameter space $\Theta$ (see [calin2014geometric]). Owing to these properties, the Fisher metric has found widespread applications in various branches of geometry (see [hitchin2006geometry, gnandi2022dual, fujiwara2020dually, groisser1997instantons, ciaglia2021focus, murray1993information, matsueda2017banados, naudts2009q, petz2007quantum, lesniewski1999monotone, facchi2010classical, barbaresco2014koszul]).
Information geometry is also deeply connected with several mathematical frameworks, including Hessian geometry, Kähler geometry, complex geometry, special Kähler geometry, and complex analysis. One of the central problems in this area concerns the realization of a statistical manifold as a statistical model, namely, a manifold endowed with dual affine connections compatible with a Fisher metric. This problem was resolved by Hông Vân Lê in [le2006statistical]. A further and significantly more difficult question is to characterize which classes of statistical manifolds can be realized as exponential families; this problem remains open. Despite its rapid progress, information geometry continues to present a number of important open questions (see [sei2025open]).
The remainder of this paper is organized as follows. Section 2 recalls the necessary background material in information geometry. Section 3 is devoted to the proof of the main result.
2. Basic Notions
This section introduces the fundamental structures studied throughout this work. For a comprehensive background and a collection of illustrative examples, we refer the reader to the following references: [amari2016information, amari2000methods, amari2012differential, amari2014curvature, ay2017information, calin2014geometric, molitor2013exponential, mehrshad2025statsitical, gnandi2022classification, gnandi2022dual, furuhata2017sasakian, molitor2025kahler, noguchi1992geometry, pal2024introduction, lauritzen1987statistical, cencov2000statistical, ay2002dually, zhang2007note, furuhata2016submanifold, matsuzoe2010statistical, takano2006statistical].
@@ 2.1. Statistical models
[amari1987differential, amari2000methods, amari2012differential, amari2014curvature, amari2016information, ay2017information, calin2014geometric] Let $\mathcal{S}$ be a family of probability distributions represented as
@@
where $\Theta$ is a domain in $\mathbb{R}^{n}$ with Cartesian coordinates $(\theta^{1},\theta^{2},\dots,\theta^{n})$, called the parameter space. Assume further that the mapping $\theta\mapsto p(\cdot,\theta)$ is bijective. Then $\mathcal{S}$ is called a statistical model determined by $p$.
@@
$\Gamma^{(\alpha)}_{ijk}(\theta)$ $=\mathbb{E}_{p(\cdot,\theta)}\Bigg[\Bigg(\frac{\partial^{2}\log p(x,\theta)}{\partial\theta^{i}\partial\theta^{j}}+\frac{1-\alpha}{2}\frac{\partial\log p(x,\theta)}{\partial\theta^{i}}\frac{\partial\log p(x,\theta)}{\partial\theta^{j}}\Bigg)$
$\qquad\qquad\cdot\log p(x,\theta)\,\frac{\partial\log p(x,\theta)}{\partial\theta^{k}}\Bigg].$
In the literature, the family of $\alpha$-connections $\nabla^{(\alpha)}$ is commonly referred to as the Amari-Chentsov $\alpha$-connections or Amari–Čencov $\alpha$-connections [amari2000methods] . This terminology honors the foundational contributions of Shun-ichi Amari [amari1987differential, amari2012differential] and Nikolai Čencov[cencov2000statistical] to information geometry. The Amari-Chentsov $\alpha$-connections provide a unified framework for studying the differential geometric structure of statistical models, with the Fisher metric as the central Riemannian structure and the $\alpha$-connections forming a dualistic structure on the statistical manifold. The $\alpha$-connection $\nabla^{(\alpha)}$ satisfies the following fundamental properties:
@@
This corresponds exactly to the Gaussian distribution $\mathcal{N}(\mu,\sigma)$. For an exponential family with natural parameters $\theta=(\theta^{1},\dots,\theta^{n})$, the Fisher metric and the cubic form are given in terms of the cumulant generating function $\psi(\theta)$ by:
@@ 2.2. Statistical manifold
Building upon the geometric insights developed by Amari and Chentsov, Steffen L. Lauritzen provided a rigorous formalization of the notion of a statistical manifold.
@@
The following construction provides a fundamental example of how statistical structures naturally arise on Riemannian manifolds equipped with a distinguished vector field. Let $(M,g)$ be a Riemannian manifold with Levi–Civita connection $\nabla^{\mathrm{LC}}$, and let $\xi\in TM\setminus\{0\}$ be a non-vanishing vector field. For all $X,Y\in\mathfrak{X}(M)$, define
@@
Then $(M,g,\nabla,\nabla^{*})$ is a statistical manifold. Starting from our previous construction, on an almost contact metric manifold $(M^{2d+1},\phi,g,\alpha,\xi)$ with $\nabla^{LC}$ the Levi–Civita connection of the metric $g$ (see [blair2010riemannian] for exposition on almost contact metric geometry), we can construct the following one-parameter families of statistical connections:
@@
For a statistical manifold $(M,g,\nabla,\nabla^{*})$, the connection $\nabla$ is flat if and only if $\nabla^{*}$ is flat. In this case, we say that $(M,g,\nabla,\nabla^{*})$ is a dually flat manifold([amari2016information]). It was shown in [ay2002dually] that any compact dually flat manifold necessarily has an infinite fundamental group. Consequently, the sphere $\mathbb{S}^{q}$, for $q>1$, cannot be a dually flat manifold.
@@ 2.3. Hessian manifold
Hessian manifolds exhibit particularly rich topological and geometric structures and share a deep relationship with Kähler manifolds. This relationship is especially manifested in the fact that the tangent bundle of any Hessian manifold is naturally endowed with a Kähler metric induced by the Hessian structure itself [shima2007geometry, dombrowski1962geometry, satoh2019almost]. Such manifolds occupy a central place in information geometry, Kähler geometry, special Kähler geometry, and complex analysis. They also play an important role in various applied fields, including convex optimization, statistics, combinatorics, Souriau’s thermodynamics on Lie groups[souriau1997, barbaresco2015koszul].
@@
where $(x_{1},\dots,x_{n})$ is a system of affine coordinates with respect to $\nabla$. A locally flat manifold $(M,\nabla)$ equipped with a Hessian metric $g$ is called a Hessian manifold and is denoted by $(M,g,\nabla)$.
@@
On an orientable Hessian manifold $(M,g,\nabla)$, using the connection $\nabla$ and the volume element induced by $g$, Koszul introduced a closed 1-form, which Shima [shima2007geometry] refers to as the first Koszul form, associated with the Hessian structure $(\nabla,g)$.
@@ 3. Main Results
Theorem 3.1.
[le2006statistical] Every smooth compact statistical manifold $(M^{n},g,T)$ of finite dimension admits, for a sufficiently large integer $N$, an isostatistical embedding (i.e., an embedding that preserves both the tensor $T$ and the metric $g$) into $\mathrm{Cap}^{N}_{+}$, the space of all positive probability distributions defined on a discrete sample space $\Omega^{N}$ consisting of $N$-elementary events, where
$\mathrm{Cap}^{N}_{+}:=\left\{(p_{1},\ldots,p_{n})\;\middle|\;p_{i}>0\text{ for }i=1,\ldots,n\;\text{and}\;\sum_{i=1}^{n}p_{i}=1\right\}.$
This result shows that compact statistical manifolds of finite dimension can indeed be realized as statistical models, although it does not provide an explicit form for such an embedding.
@@
On any compact statistical manifold, there exists a singular foliation whose leaves are Hessian manifolds. Every compact, orientable leaf is either a finite quotient of a flat torus or a mapping torus with a periodic gluing map. If such a leaf is three-dimensional and does not admit any flat metric, then all of its Betti numbers are odd, and it arises as a quotient of an exponential family.
@@
The proof proceeds in four steps. The first step is presented in the PhD thesis([gnandi2022classification]), but we have deemed it useful to provide additional details to facilitate understanding for the readers of this paper.
@@ 3. Step 3:
    We thus deduce that $\widetilde{Z}$ is the Reeb vector field associated with the cosymplectic structure $(\widetilde{\beta},\Omega)$ (see [li2008topology]). Since $\widetilde{Z}$ is a Killing vector field on $L_{\nabla}$, it follows from [bazzoni2015k] that $(L_{\nabla},\widetilde{\beta},\Omega)$ is a K-cosymplectic manifold. By [bazzoni2015k, Proposition 2.8], $L_{\nabla}$ admits an K-cosymplectic structure $(\widetilde{\beta},\widetilde{Z},\phi,h)$ with $\mathcal{L}_{\widetilde{Z}}h=0$.
@@
    $L_{\nabla}\cong\frac{\mathbb{H}^{2}\times\mathbb{R}}{\Gamma}\quad\text{with}\quad\Gamma\subset\mathrm{Isom}_{0}(\mathbb{H}^{2}\times\mathbb{R})\cong\mathrm{Isom}_{0}(\mathbb{H}^{2})\times\mathrm{Isom}_{0}(\mathbb{R}).$
@@ 4. Step 4:
    (3.1) $P((x,y,z),(\xi,\eta,\zeta))=\exp\left(-(x\xi+y\eta+z\zeta)+\frac{3}{2}\log q-\log\chi(0,0,1)\right).$
@@
    Introducing cylindrical coordinates $(r,\alpha,t)$ by
@@
    Here $(r,\alpha)$ are polar coordinates on $\mathbb{H}^{2}$ and $t$ is the coordinate $\mathbb{R}$. Thus, $(\mathbb{H}^{2}\times\mathbb{R},g)$ is isometric to $(\Theta,g^{F})$. We deduce that, the leaf $L_{\nabla}$ of the associated foliation $\mathcal{F}_{\nabla}$ satisfies
    $L_{\nabla}\cong\frac{\Theta}{\Lambda},\qquad\Lambda\subset\mathrm{Isom}_{0}(\Theta).$
@@
    $L_{\nabla}\cong\frac{\{P_{\theta}:\theta\in\Theta\}}{\Lambda}\,\qquad\Lambda\subset\mathrm{Isom}_{0}(\{P_{\theta}:\theta\in\Theta\}),$
    where $\{P_{\theta}:\theta\in\Theta\}$ is an exponential family defined in 3.1. This completes the proof.
@@
$L_{\nabla}\cong\frac{\mathbb{H}^{2}\times\mathbb{R}}{\Gamma},\qquad\Gamma\subset\mathrm{Isom}_{0}(\mathbb{H}^{2}\times\mathbb{R}),$
is precisely the Seifert manifold with Euler number 0 and hyperbolic orbifold base $B$ (since $\chi_{\mathrm{orb}}(B)<0$). In other words, $L_{\nabla}$ admits a $\mathbb{H}^{2}\times\mathbb{R}$-geometry. From [martelli2016introduction, Prop. 10.3.26], $L_{\nabla}$ is covered by $\Sigma_{g}\times\mathbb{S}^{1}$ for some $g>1$. Hence
$L_{\nabla}\cong\frac{\Sigma_{g}\times\mathbb{S}^{1}}{\Pi},$
where $\Pi$ is a finite group acting freely and properly on $\Sigma_{g}\times\mathbb{S}^{1}$. Since $\Sigma_{g}$ admits an $\mathbb{RP}^{2}$-structure (see [benzecri1960varietes, p. 243] and [goldman2022geometric, Cor. 6.5.2]), the universal cover of $\Sigma_{g}\times\mathbb{S}^{1}$, namely $\mathbb{H}^{2}\times\mathbb{R}$, develops into the Lorentz cone $\Theta$. Consequently,
$\Sigma_{g}\times\mathbb{S}^{1}\cong\frac{\Theta}{\Gamma\times\langle\lambda\rangle},\qquad\Gamma\subset\mathrm{SO}(2,1),\quad\lambda>1.$
@@
Finally, by using the Koszul-Vinberg characteristic function $\chi$ on the sharp convex cone $\Theta$, and arguing as above, the result follows.
Discussion 3.4.
In this work, we focused primarily on the $3$-dimensional leaves of singular foliations, which we showed can be realized as quotients of exponential families under certain conditions. A natural open question remains: what occurs for leaves of higher dimensions? A deeper understanding of their geometric and topological structure could open new avenues in information geometry and statistics. Furthermore, we have mainly considered classical exponential families; extending our approach to other classes of statistical models represents a promising direction for future research.